# MeshPad: Interactive Sketch-Conditioned Artist-Designed Mesh Generation and Editing

### v2

@@ Abstract
We introduce MeshPad, a generative approach that creates 3D meshes from sketch inputs. Building on recent advances in artist-designed triangle mesh generation, our approach addresses the need for interactive mesh creation. To this end, we focus on enabling consistent edits by decomposing editing into ‘deletion’ of regions of a mesh, followed by ‘addition’ of new mesh geometry. Both operations are invoked by simple user edits of a sketch image, facilitating an iterative content creation process and enabling the construction of complex 3D meshes. Our approach is based on a triangle sequence-based mesh representation, exploiting a large Transformer model for mesh triangle addition and deletion. In order to perform edits interactively, we introduce a vertex-aligned speculative prediction strategy on top of our additive mesh generator. This speculator predicts multiple output tokens corresponding to a vertex, thus significantly reducing the computational cost of inference and accelerating the editing process, making it possible to execute each editing step in only a few seconds. Comprehensive experiments demonstrate that MeshPad outperforms state-of-the-art sketch-conditioned mesh generation methods, achieving more than $22\%$ mesh quality improvement in Chamfer distance, and being preferred by $90\%$ of participants in perceptual evaluations.
Figure 1: MeshPad enables interactive mesh creation and editing with sketches. We decompose this complex task into two sketch-conditioned operations: addition and deletion. Top: our method allows for a user to create and modify artist-designed triangle meshes by simply drawing and editing 2D sketches, achieving intuitive and interactive 3D modeling. Bottom (left): our method generates a variety of complex yet compact meshes. Bottom (right): our interactive user interface allows users to iteratively edit the mesh, with each edit step taking a few seconds.
@@ 1 Introduction
Algorithmically, MeshPad leverages a hybrid approach combining an autoregressive network for addition and a token-classification network for deletion. Unlike existing artist-designed mesh generation methods, our addition network generates partial shapes only corresponding to new sketch strokes, rather than repeatedly synthesizing the entire mesh. This naturally solves the problem of preserving unedited mesh regions during editing. Additionally, based on the triangle sequence representation [17], we introduce a vertex-aligned speculative decoder to accelerate autoregressive generation and reduce computational time. Instead of predicting one vertex coordinate at a time, the speculator allows us to predict $3$ coordinates, i.e., one vertex, at a time. We find that jointly training the speculative head and aligning it with the vertex tokens reduces generation time while maintaining generation quality. We further demonstrate our method through a user interface (Fig. 1) that allows users to interactively edit 3D meshes by drawing on 2D sketches. Extensive experiments show that our method outperforms state-of-the-art methods in sketch-conditioned mesh generation, producing cleaner meshes that more accurately align with the sketch input and excelling in partial mesh editing tasks. Finally, our speculative prediction acceleration enables interactive mesh generation and editing within seconds while maintaining overall shape quality.
@@
  • Our vertex-aligned speculative classification head for addition notably accelerates the triangle mesh generation by $\times2$ without quality loss.
@@ 2 Related Work
However, such works struggle to produce an artist-designed mesh. Techniques based on polygonal meshes or other explicit representations can rely on an expensive optimization [2, 24], deformation priors [54], specialized parametric families [22] or procedural generators [62]. Such approaches do not generalize easily to arbitrary geometry and have difficulties making local edits without conflicting with the global constraints. In contrast, our approach addresses these issues using a sequence-based mesh representation to perform partial editing by removing and adding triangles.
@@ 3.1 Mesh Creation and Editing
We decouple sketch-based mesh creation and editing into two subtasks: addition and deletion. Both operations are conditioned on a sketch image $\mathcal{I}$, a bitmap where colored line strokes indicate regions to be added or deleted. We define a mesh $\mathcal{M}$ as a set of triangles ${}\mathcal{M}=\{\mathcal{F}\}$, where triangles =$\mathcal{F}=\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}\}$ and $\boldsymbol{v}\in\mathbb{R}^{3}$ represents vertices.
@@
Figure 3: Model Architecture. We use the Open Pre-trained Transformer (OPT) as the backbone to process the embedded triangle sequence and the sketch embeddings from a pre-trained image foundation model, RADIO (frozen during training). Top (right): the deletion network is a classification network that labels each input mesh vertex for removal. The deletion head takes the $xyz$ coordinate embeddings of a vertex as input and produces a masking label for the vertex. Bottom (right): the addition network is an autoregressive generation model that predicts new mesh tokens with a speculator. During the speculative prediction, we align the speculator to the vertex axis so that it always takes the $x$ coordinate as input and predicts the vertex’s $y$ and $z$ coordinates. E, M, V, and <s> represent Transformer hidden states, vertex masks, vertex coordinate tokens, and the <split> token, respectively.
@@ 3.2 Network Architecture
We detokenize the generated sequence $S^{\prime}_{r}$ to mesh triangles and then merge them with the input partial mesh:
@@ 3.4 Training Data Generation for Mesh Editing
=$\mathcal{M}=\left\{\mathcal{F}\in\mathcal{M}_{c}|\exists\boldsymbol{p}\in\mathcal{F}:\boldsymbol{p}\in\mathcal{L}\right\},$ (6)
@@
Figure 4: An example of training data generation: the volume $\mathcal{L}$ is sampled to cover a large portion from the top, while $\mathcal{L}_{k}$ covers the chair backrest. The mesh $\mathcal{M}$ and $\mathcal{M}_{k}$ contains triangles with any vertex in $\mathcal{L}$ and $\mathcal{L}_{k}$, respectively. $\mathcal{M}_{r}$ is the difference between $\mathcal{M}$ and $\mathcal{M}_{k}$. The red and black line strokes in sketch $\mathcal{I}$ corresponds to $\mathcal{M}_{k}$, and $\mathcal{M}_{r}$, respectively.
@@ 4.1 Implementation Details
Data. Following MeshGPT [51] and MeshAnythingV2 [17], we preprocess and filter the ShapeNet dataset [9] to obtain approximately $28$k meshes from $55$ categories, which each have a face count $<768$ for training. We sample around $500$ meshes for validation and $1000$ meshes for testing. Other meshes are used for training.
@@ 4.2 Metrics and Baselines
We further conduct both unary and binary perceptual studies. In the unary study, participants evaluate mesh quality (GQ) and sketch matching (GM) for mesh generation. For mesh editing, they rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part. Each is rated on a scale of $1$ to $5$ ($=$best). In the binary study, we ask participants to compare our method with baselines and select their preferred method. We report the percentages of preferences for our method compared to the baselines.
@@
Figure 6: Sequence visualization of mesh creation with MeshPad. We show the input sketch on the top left of each of our method outputs. Green and purple arrows represent addition and deletion, respectively. The “$+$” inside a green arrow indicates multiple addition steps.
@@ 4.3 Experimental Setup
We use the test set described in Sec. 4.1 and automatically generate sketches from one of five views: top-left, top-front-left, top-front, top-front-right, or top-right, for the generation task. To further show the robustness of our method, we use the IKEA dataset [36, 53], which contains $188$ furniture models. We use automatically generated sketches (follow Sec. 3.4) for generation tasks because it helps to analyze the sketch-to-mesh correspondence by the CLIP and LPIPS scores. In our qualitative experiments we show our method generalizes to hand-drawn sketches. Additionally, we randomly select $50$ shapes of airplanes, chairs, and lamps (on which SENS is trained) from the two datasets, automatically generate their sketches for generation, and then manually edit the sketches for one-step editing evaluation.
For the generation task, we perform one-step addition operation for our method. We provide additional camera view conditions in all LAS runs as suggested in [63]. We also apply dilation and resizing to sketches to match the input to the training data of LAS. As SENS is category-specific, we use its checkpoint corresponding to the category of the target mesh for each evaluation run.
@@ 4.4 Comparison to State of the Art
Table 2: Sketch-based metrics and unary perceptual study ratings (ranged $1$-$5$) on generation and editing results of our hand-drawn sketch evaluation set. We benchmark against LAS [63] and SENS [7], with outputs further refined using MeshAnythingV2 (MA) [17]. For mesh generation, participants evaluate mesh quality (GQ) and sketch matching (GM). For mesh editing, participants rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part.
@@ 4.5 Ablations
Vertex-alignment preserves mesh quality while achieving comparable speedup as vanilla MLP speculator. Comparing no speculator (w/o speculator) and speculator without vertex alignment (w/o vert-alignment), we can see that the speculator, without aligning to vertex tokens, results in a notable speedup, but at large penalty cost in mesh quality and sketch-to-mesh correspondence. By aligning the speculator to vertex tokens, our method reaches the same level of sketch-to-mesh correspondence as the no speculator version while maintaining the naive speculator’s advantage in generation speed. Additionally, applying the vertex-aligned speculator results in slightly higher metrics in mesh quality. This is likely due to the vertex-aligned speculator reducing the complexity of the linear prediction head. Rather than predicting control tokens and all $x$, $y$, and $z$ coordinate tokens, it is trained to predict only control tokens and $x$ coordinate tokens, which reduces the overall prediction error. Thus, equipped with the vertex-aligned speculator, our method achieves strong performance and fast generation speeds (approximately $26.7$ faces per second considering the average face density of the mesh sequences). As a result, mesh additions, which typically generate about $\nicefrac{{1}}{{5}}$ of a complete shape, run in about 1 to 5 seconds (depending on the number of faces added), enabling interactive runtimes for editing.
@@
Table 4: Ablation study. We ablate our speculator, vertex alignment of our speculator (using a common MLP speculator [57] with the same length of $2$), and its joint training with our mesh transformer. The unit of CD is 0.001. T/s represents tokens per second and is measured on NVIDIA A100. Our vertex-aligned speculator accelerates generation by 2.16×2.16$\times$2.16 × without loss in mesh quality. Compared to an MLP speculator without vertex-alignment, which suffers from quality loss, our method runs at a comparable speed.
@@ Acknowledgements
In this material, we first introduce the details of sequentialized mesh representation in Appendix A, followed by details of our self-supervised data processing in Appendix B. We introduce the implementation of our user interface in Appendix C. Then, we showcase three additional qualitative experiments in Appendix D. Next, we provide more visual evaluation results in Appendix E followed by more details and results from our perceptual study in Appendix F.
@@ Appendix A Mesh Token Sequence Details
As stated in Sec. 3.2, The sequential representation of a mesh consists of vertex coordinate tokens and control tokens where the <split> token separates the full sequence representing the whole mesh into subsequences corresponding to a “triangle chain” of adjacent triangles. Each subsequence has $3n$ vertex coordinate tokens where $n$ is the number of vertices contained in the subsequence. Given a subsequence, we first decode a sequence of vertices $\{\boldsymbol{v}_{k}\}_{k=0}^{n}$: the $x$, $y$, and $z$ position of $\boldsymbol{v}_{k}$ is obtained from the coordinate tokens at positions $3k$, $13k+$ 1, and $23k+$ 2 in the subsequence, respectively. We then get the triangle chain by forming triangles of three adjacent vertices in the vertex sequence:
{}=$\{\mathcal{F}\}=\{\{\boldsymbol{v}_{k},\boldsymbol{v}_{k-1},\boldsymbol{v}_{k-2}\}|\forall k:k>=2\}.$ (7)
Merging all triangle chains obtained from all subsequences results in the full mesh.
@@ Appendix B Data Generation Details
Random sampling of volumes. To sample volumes that satisfy $\mathcal{L}_{k}\subsetneq\mathcal{L}$, we first randomly define two volumes $\mathcal{L}_{\{a,b\}}$ within the bounding volume (assuming a unit cube) of the complete mesh $\mathcal{M}_{c}$. For each volume $\tilde{\mathcal{L}}$ of $\mathcal{L}_{\{a,b\}}$, we sample the volume as following: first, randomly choose an axis $i$ from $\{1,2,3\}$ indicating the $x$, $y$, or $z$ axis, then randomly select one region $\mathcal{R}$ from the four candidates: $[-\infty,a]$, $[a,+\infty]$, $[b-c,b+c]$, $[-\infty,b-c]\cup[b+c,+\infty]$, where $a\in[0.2,0.8]$, $b\in[0.4,0.6]$ and $c\in[0.1,0.4]$ are uniformly sampled within their defined range. The sampled volume is then defined as:
~=$\tilde{\mathcal{L}}=\{\boldsymbol{p}\in\mathbb{R}^{3}|p_{i}\in\mathcal{R}\}.$ (8)
@@
Sketch generation. Automatic sketch generation is done by applying Canny edge detection on the rendered depth and normal images of $\mathcal{M}$. We then merge the two edge detection results to obtain the corresponding synthetic sketch used for training. To get mutually exclusive sets of line strokes $\mathcal{I}_{\{r,k\}}$, which correspond to different parts of the mesh, we further render a visibility mask of $\mathcal{M}_{r}$ in the same camera view and collect all line strokes within the mask as $\mathcal{I}_{r}$. $\mathcal{I}_{k}$ is then defined as the line strokes not in $\mathcal{I}_{r}$. In practice, the line strokes from the Canny edge detector could lie outside the mask by $1$ pixel. Therefore, we dilate the visibility mask by $1$ pixel before collecting $\mathcal{I}_{r}$. During training, we randomly sample camera views from the upper unit viewing hemisphere, with azimuth angles in $[-90^{\circ},90^{\circ}]$ and elevation angles in $[0^{\circ},60^{\circ}]$.

### v3

# MeshPad: Interactive Sketch-Conditioned Artist-Reminiscent Mesh Generation and Editing

@@ Abstract
We introduce MeshPad, a generative approach that creates 3D meshes from sketch inputs. Building on recent advances in artist-reminiscent triangle mesh generation, our approach addresses the need for interactive mesh creation. To this end, we focus on enabling consistent edits by decomposing editing into ‘deletion’ of regions of a mesh, followed by ‘addition’ of new mesh geometry. Both operations are invoked by simple user edits of a sketch image, facilitating an iterative content creation process and enabling the construction of complex 3D meshes. Our approach is based on a triangle sequence-based mesh representation, exploiting a large Transformer model for mesh triangle addition and deletion. In order to perform edits interactively, we introduce a vertex-aligned speculative prediction strategy on top of our additive mesh generator. This speculator predicts multiple output tokens corresponding to a vertex, thus significantly reducing the computational cost of inference and accelerating the editing process, making it possible to execute each editing step in only a few seconds. Comprehensive experiments demonstrate that MeshPad outperforms state-of-the-art sketch-conditioned mesh generation methods, achieving more than $22\%$ mesh quality improvement in Chamfer distance, and being preferred by $90\%$ of participants in perceptual evaluations.
Figure 1: MeshPad enables interactive mesh creation and editing with sketches. We decompose this complex task into two sketch-conditioned operations: addition and deletion. Top: our method allows for a user to create and modify artist-reminiscent triangle meshes by simply drawing and editing 2D sketches, achieving intuitive and interactive 3D modeling. Bottom (left): our method generates a variety of complex yet compact meshes. Bottom (right): our interactive user interface allows users to iteratively edit the mesh, with each edit step taking a few seconds.
@@ 1 Introduction
Algorithmically, MeshPad leverages a hybrid approach combining an autoregressive network for addition and a token-classification network for deletion. Unlike existing artist-reminiscent mesh generation methods, our addition network generates partial shapes only corresponding to new sketch strokes, rather than repeatedly synthesizing the entire mesh. This naturally solves the problem of preserving unedited mesh regions during editing. Additionally, based on the triangle sequence representation [17], we introduce a vertex-aligned speculative decoder to accelerate autoregressive generation and reduce computational time. Instead of predicting one vertex coordinate at a time, the speculator allows us to predict $3$ coordinates, i.e., one vertex, at a time. We find that jointly training the speculative head and aligning it with the vertex tokens reduces generation time while maintaining generation quality. We further demonstrate our method through a user interface (Fig. 1) that allows users to interactively edit 3D meshes by drawing on 2D sketches. Extensive experiments show that our method outperforms state-of-the-art methods in sketch-conditioned mesh generation, producing cleaner meshes that more accurately align with the sketch input and excelling in partial mesh editing tasks. Finally, our speculative prediction acceleration enables interactive mesh generation and editing within seconds while maintaining overall shape quality.
@@
  • Our vertex-aligned speculative classification head for addition notably accelerates the triangle mesh generation by $2.2\times$ without quality loss.
@@ 2 Related Work
However, such works struggle to produce an artist-reminiscent mesh. Techniques based on polygonal meshes or other explicit representations can rely on an expensive optimization [2, 25], deformation priors [56], specialized parametric families [23] or procedural generators [64]. Such approaches do not generalize easily to arbitrary geometry and have difficulties making local edits without conflicting with the global constraints. In contrast, our approach addresses these issues using a sequence-based mesh representation to perform partial editing by removing and adding triangles.
@@ 3.1 Mesh Creation and Editing
We decouple sketch-based mesh creation and editing into two subtasks: addition and deletion. Both operations are conditioned on a sketch image $\mathcal{I}$, a bitmap where colored line strokes indicate regions to be added or deleted. We define a mesh $\mathcal{M}$ as a set of triangles $\mathcal{M}=\{\mathcal{F}\}$, where triangles $\mathcal{F}=\{\boldsymbol{v}_{1},\boldsymbol{v}_{2},\boldsymbol{v}_{3}\}$ and $\boldsymbol{v}\in\mathbb{R}^{3}$ represents vertices.
@@
Figure 3: Model Architecture. We use the Open Pre-trained Transformer (OPT) in both the addition and deletion networks as the backbone to process the embedded triangle sequence and the sketch embeddings from a pre-trained image foundation model, RADIO (frozen during training). Top (right): the deletion network is a classification network that labels each input mesh vertex for removal. The deletion head takes the xyzxyz coordinate embeddings of a vertex as input and produces a masking label for the vertex. Bottom (right): the addition network is an autoregressive generation model that predicts new mesh tokens with a speculator. The CTransformer backbone uses causal attention instead of full attention. During the speculative prediction, we align the speculator to the vertex axis so that it always takes the $x$ coordinate as input and predicts the vertex’s $y$ and $z$ coordinates. E, M, V, and <s> represent Transformer hidden states, vertex masks, vertex coordinate tokens, and the <split> token, respectively.
@@ 3.2 Network Architecture
The token $S_{r}^{\prime(i+1)}$ is obtained by sampling with the distribution $P$, and then conditions the generation of $S_{r}^{\prime(i+2)}$. We detokenize the generated sequence $S^{\prime}_{r}$ to mesh triangles and then merge them with the input partial mesh:
@@ 3.4 Training Data Generation for Mesh Editing
$\mathcal{M}=\left\{\mathcal{F}\in\mathcal{M}_{c}|\exists\boldsymbol{p}\in\mathcal{F}:\boldsymbol{p}\in\mathcal{L}\right\},$ (6)
@@
Figure 4: An example of training data generation: the volume $\mathcal{L}$ is sampled to cover a large portion from the top, while $\mathcal{L}_{k}$ covers the chair backrest. The meshes $\mathcal{M}$ and $\mathcal{M}_{k}$ contain triangles with any vertex in $\mathcal{L}$ and $\mathcal{L}_{k}$, respectively—that is, triangles fully or partially inside the volumes. $\mathcal{M}_{r}$ is the difference between $\mathcal{M}$ and $\mathcal{M}_{k}$. The red and black line strokes in sketch $\mathcal{I}$ corresponds to $\mathcal{M}_{k}$, and $\mathcal{M}_{r}$, respectively.
@@ 4.1 Implementation Details
Data. Following MeshGPT [53] and MeshAnythingV2 [17], we preprocess and filter the ShapeNet dataset [9] to obtain approximately 2828k meshes from $55$ categories, which each have a face count <768<768 for training. We sample around $500$ meshes for validation and $1000$ meshes for testing. Other meshes are used for training.
@@ 4.2 Metrics and Baselines
We further conduct both unary and binary perceptual studies. In the unary study, participants evaluate mesh quality (GQ) and sketch matching (GM) for mesh generation. For mesh editing, they rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part. Each is rated on a scale of $1$ to 55 (==best). In the binary study, we ask participants to compare our method with baselines and select their preferred method. We report the percentages of preferences for our method compared to the baselines.
@@
Figure 6: Sequence visualization of mesh creation with MeshPad. We show the input sketch on the top left of each of our method outputs. Green and purple arrows represent addition and deletion, respectively. The “+++” inside a green arrow indicates multiple addition steps.
@@ 4.3 Experimental Setup
We use the test set described in Sec. 4.1 and automatically generate sketches from one of five views: top-left, top-front-left, top-front, top-front-right, or top-right, for the generation task. To further show the robustness of our method, we use the IKEA dataset [38, 55], which contains $188$ furniture models. We use automatically generated sketches (following Sec. 3.4) for generation tasks because they aid the analysis of sketch-to-mesh correspondence by the CLIP and LPIPS scores. In our qualitative experiments, we show that our method generalizes to hand-drawn sketches. Additionally, we randomly select $50$ shapes of airplanes, chairs, and lamps (on which SENS is trained) from the two datasets, automatically generate their sketches for generation, and then manually edit the sketches for one-step editing evaluation.
For the generation task, our method predicts an entire mesh from scratch conditioned on the sketch input. We provide additional camera view conditions in all LAS runs as suggested in [65]. We also apply dilation and resizing to sketches to match the input to the training data of LAS. As SENS is category-specific, we use its checkpoint corresponding to the category of the target mesh for each evaluation run.
@@ 4.4 Comparison to State of the Art
Table 2: Sketch-based metrics and unary perceptual study ratings (ranged 11-55) on generation and editing results of our hand-drawn sketch evaluation set. We benchmark against LAS [65] and SENS [7], with outputs further refined using MeshAnythingV2 (MA) [17]. For mesh generation, participants evaluate mesh quality (GQ) and sketch matching (GM). For mesh editing, participants rate edited mesh quality (EQ), edited sketch matching (EM), and edited mesh consistency (EC) of the unedited part.
@@ 4.5 Ablations
Vertex-alignment preserves mesh quality while achieving comparable speedup as vanilla MLP speculators. Comparing no speculator (w/o speculator) and speculator without vertex alignment (w/o vert-alignment), we can see that the speculator, without aligning to vertex tokens, results in a notable speedup, but at large penalty cost in mesh quality and sketch-to-mesh correspondence. By aligning the speculator to vertex tokens, our method reaches the same level of sketch-to-mesh correspondence as the no speculator version while maintaining the naive speculator’s advantage in generation speed. Additionally, applying the vertex-aligned speculator results in slightly higher metrics in mesh quality. This is likely due to the vertex-aligned speculator reducing the complexity of the linear prediction head. Rather than predicting control tokens and all $x$, $y$, and $z$ coordinate tokens, it is trained to predict only control tokens and $x$ coordinate tokens, which reduces the overall prediction error. Thus, equipped with the vertex-aligned speculator, our method achieves strong performance and fast generation speeds (approximately $26.7$ faces per second considering the average face density of the mesh sequences). As a result, mesh additions, which typically generate about $\nicefrac{{1}}{{5}}$ of a complete shape, run in about 1 to 5 seconds (depending on the number of faces added), enabling interactive runtimes for editing.
@@
Table 4: Ablation study. We ablate our speculator, vertex alignment of our speculator (using a common MLP speculator [59] with the same length of $2$), and its joint training with our mesh transformer. The unit of CD is 0.001. T/s represents tokens per second and is measured on NVIDIA A100. Our vertex-aligned speculator accelerates generation by $2.16\times$ without loss in mesh quality. Compared to an MLP speculator without vertex-alignment, which suffers from quality loss, our method runs at a comparable speed.
@@ Acknowledgements
In this material, we first introduce the details of sequentialized mesh representation in Appendix A, followed by details of our self-supervised data processing in Appendix B. We introduce the implementation of our user interface in Appendix C. Then, we showcase four additional qualitative experiments in Appendix D. Next, we provide more visual evaluation results in Appendix E and topological statistics of generated meshes in Appendix F, followed by more details and results from our perceptual study in Appendix G.
@@ Appendix A Mesh Token Sequence Details
As stated in Sec. 3.2, The sequential representation of a mesh consists of vertex coordinate tokens and control tokens where the <split> token separates the full sequence representing the whole mesh into subsequences corresponding to a “triangle chain” of adjacent triangles. Each subsequence has $3n$ vertex coordinate tokens where $n$ is the number of vertices contained in the subsequence. Given a subsequence, we first decode a sequence of vertices $\{\boldsymbol{v}_{k}\}_{k=0}^{n}$: the $x$, $y$, and $z$ position of $\boldsymbol{v}_{k}$ is obtained from the coordinate tokens at positions $3k$, $3k+1$, and $3k+2$ in the subsequence, respectively. We then get the triangle chain by forming triangles of three adjacent vertices in the vertex sequence:
$\{\mathcal{F}\}=\{\{\boldsymbol{v}_{k},\boldsymbol{v}_{k-1},\boldsymbol{v}_{k-2}\}|\forall k:k>=2\}.$ (7)
Merging all triangle chains obtained from all subsequences results in the full mesh. When converting a mesh into sequence, we first sort every triangle in the order of zyxz-y-x, then perform a depth-first traversal from the first triangle to find adjacent triangles as the first subsequence. We then subtract the triangles already traversed and start from a new triangle to obtain the next subsequence. We iterate this process until all triangles are traversed.
@@ Appendix B Data Generation Details
Random sampling of volumes. To sample volumes that satisfy $\mathcal{L}_{k}\subsetneq\mathcal{L}$, we first randomly define two volumes $\mathcal{L}_{\{a,b\}}$ within the bounding volume (assuming a unit cube) of the complete mesh $\mathcal{M}_{c}$. For each volume $\tilde{\mathcal{L}}$ of $\mathcal{L}_{\{a,b\}}$, we sample the volume as following: first, randomly choose an axis $i$ from $\{1,2,3\}$ indicating the $x$, $y$, or $z$ axis, then randomly select one region $\mathcal{R}$ from the four candidates: $\infty,a]$, $a,+\infty]$, [bc,b+c][b-c,b+c], $\infty,b-c]\cup[b+c,+\infty]$, where $a\in[0.2,0.8]$, $b\in[0.4,0.6]$ and $c\in[0.1,0.4]$ are uniformly sampled within their defined range. The sampled volume is then defined as:
$\tilde{\mathcal{L}}=\{\boldsymbol{p}\in\mathbb{R}^{3}|p_{i}\in\mathcal{R}\}.$ (8)
@@
Sketch generation. Automatic sketch generation is done by applying Canny edge detection on the rendered depth and normal images of $\mathcal{M}$. We then merge the two edge detection results to obtain the corresponding synthetic sketch used for training. To get mutually exclusive sets of line strokes $\mathcal{I}_{\{r,k\}}$, which correspond to different parts of the mesh, we further render a visibility mask of $\mathcal{M}_{r}$ in the same camera view and collect all line strokes within the mask as $\mathcal{I}_{r}$. $\mathcal{I}_{k}$ is then defined as the line strokes not in $\mathcal{I}_{r}$. In practice, the line strokes from the Canny edge detector could lie outside the mask by $1$ pixel. Therefore, we dilate the visibility mask by $1$ pixel before collecting $\mathcal{I}_{r}$. During training, we randomly sample camera views from the upper unit viewing hemisphere, with azimuth angles in $90^{\circ},90^{\circ}]$ and elevation angles in $0^{\circ},60^{\circ}]$.